# HOI-Diff: Text-Driven Synthesis of 3D Human-Object Interactions using Diffusion Models

### 11 Dec 2023 v1

@@ Abstract
We address the problem of generating realistic 3D human-object interactions (HOIs) driven by textual prompts. Instead of a single model, our key insight is to take a modular design and decompose the complex task into simpler sub-tasks. We first develop a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioning on the input text, and encourage coherent motions by a cross-attention communication module between the human and object motion generation branches. We also develop an affordance prediction diffusion model (APDM) to predict the contacting area between the human and object during the interactions driven by the textual prompt. The APDM is independent of the results by the HOI-DM and thus can correct potential errors by the latter. Moreover, it stochastically generates the contacting points to diversify the generated motions. Finally, we incorporate the estimated contacting points into the classifier-guidance to achieve accurate and close contact between humans and objects. To train and evaluate our approach, we annotate BEHAVE dataset with text descriptions. Experimental results demonstrate that our approach is able to produce realistic HOIs with various interactions and different types of objects.
Figure 1: HOI-Diff can generate realistic motions for 3D human-object interactions given a text prompt and object geometry. Please see the supplementary material for video results. Darker color indicates later frames in the sequence. Best viewed in color.
@@ 1 Introduction
The generation of natural and physically plausible 3D HOIs involves humans interacting with dynamic objects in various ways according to the text prompts, thereby posing several challenges. First, the variability of object shapes makes it particularly challenging to generate semantically meaningful contact between the human and objects to avoid floating objects. Second, the generated HOIs should be faithful to the input text prompts as there are many plausible interactions between human and the same object (e.g., a person carries a chair, sits on a chair, pushes or pulls a chair). Text-driven 3D HOI synthesis with a diverse set of interactions is still under-explored. Third, the development and evaluation of 3D HOI synthesis models requires a high-quality human motion dataset with various HOIs and textual descriptions, but existing datasets lack either HOIs [Guo_2022_CVPR, kit2016] or textual descriptions [bhatnagar22behave].
@@
On one hand, recent methods [kulkarni2023nifty, jiang2022chairs, hassan_samp_2021, starke2019neural, zhang2022couch, wu2022saga, taheri2021goal, Pi_2023_ICCV] can synthesize realistic human motions for HOIs for static objects only. They usually synthesize the motion in the last mile of interaction, i.e. the motion between the given starting human pose and the final interaction pose, and overlook the movement of the objects when the human is interacting with them. On the other hand, existing methods for motion generation with dynamic objects do not adequately reflect the real-world complexity. For instance, they focus on grasping small objects [ghosh2023imos], provide the object motion as conditioning [li2023object], predict deterministic interactions between the human and the same object without the diversity [xu2023interdiff, razali2023action], consider only a small set of interactions (e.g. sit/lift [kulkarni2023nifty], sit/lie down [hassan_samp_2021], sit [jiang2022chairs, zhang2022couch, Pi_2023_ICCV], grasp [wu2022saga, taheri2021goal]), or investigate a single type of object (e.g. chair [jiang2022chairs, zhang2022couch]).
@@
In this paper, we go beyond and propose HOI-Diff for 3D HOIs synthesis involving humans interacting with different types of objects in diverse ways, which are both physically plausible and semantically faithful to the textual prompt, as shown in Fig. 1. Our key insight is to decompose 3D HOIs synthesis into three modules: (a) coarse 3D HOIs generation that extends the human motion diffusion model [tevet2023human] to a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioning on the input text prompt. To encourage coherent motions, we develop a cross-attention communication module, exchanging information between the human and object motion generation models; (b) affordance prediction diffusion model (APDM) that estimates the contacting points between the human and object during the interactions driven by the textual prompt. Our APDM does not rely on the results of the HOI-DM and thus can recover from its potential errors. Moreover, it stochastically generates the contacting points to diversity the generated motions; and (c) affordance-guided interaction correction that incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects. Compared with designing a monolithic model, our proposed HOI-Diff disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of three modules.
@@
To our best knowledge, HOI-Diff is the first text-driven approach capable of generating realistic motions for various HOIs and different types of objects during the interactions. For both training and evaluation purposes, we annotate each video sequence in BEHAVE dataset [bhatnagar22behave] with text descriptions, which bridges the gap of no data for text-driven 3D HOIs generation. Extensive experiments validate the effectiveness and design choices of our approach, particularly for dynamic objects, thereby enabling a set of new applications in human motion generation.
@@ 2 Related Work
Human Motion Generation with Diffusion Models. The denoising diffusion models have been widely used 2D image generations [rombach2022high, saharia2022photorealistic, ramesh2021zero] and achieved impressive results. Recent work [zhang2022motiondiffuse, tevet2023human, chen2023executing, karunratanakul2023gmd, rempeluo2023tracepace, ahn2023can, barquero2023belfusion, chen2023humanmac, dabral2023mofusion, shafir2023human, sun2023towards, tian2023transfusion, wei2023understanding, zhang2023generating, zhang2023remodiffuse, zhang2023tedi, xie2023omnicontrol] apply the diffusion model in the task of human motion generation. While these methods have successfully generated human motion, they usually generate isolated motions in the free space without considering the objects the human is interacting with. Our method is primarily focused on motion generation with human-object interactions.
@@
Scene- and Object-Aware Human Motion Generation. Recent works condition motion synthesis on scene geometry [huang2023diffusion, zhao2023synthesizing, wang2022towards, wang2022humanise]. This facilitates the understanding of human-scene interactions. However, the motion fidelity is compromised due to the lack of paired full scene-motion data. Other approaches [kulkarni2023nifty, jiang2022chairs, hassan_samp_2021, starke2019neural, zhang2022couch, Pi_2023_ICCV] instead focus on the interactions with the objects and can produce realistic motions. However, they focus on interacting with static objects with limited interactions. OMOMO [li2023object] proposes a conditional diffusion framework that can generate full-body motion from the object motion. The object motion is needed as input in OMOMO, whereas our method can jointly synthesize human motion and object motion. IMoS [ghosh2023imos] synthesizes the full-body human along with the 3D object motions from textual inputs. Although they can generate realistic interactions, IMoS only focuses on grasping small objects with hands. InterDiff [xu2023interdiff] predicts whole-body interactions with dynamic objects. However, it focuses on the motion prediction task, where the future motion needs to be forecasted based on past motion information and object information. Note that the interaction type is deterministic, given the object and past motion information. Different from this, we tackle the motion synthesis task, where the interaction with the same object can be controlled by the text prompt.
@@
Multi-Human Motion Generation. Recent works [shafir2023human, liang2023intergen, shafir2023human] incorporate human-to-human interactions into the motion diffusion process. They usually also adopt cross-attention to exchange information from different human motions. Different from them, our goal is to generate coarse HOIs instead of doing so in a single run, which alleviates the burden of the model.
@@
Affordance Estimation. The affordance estimation on 3D point cloud is studied in [ngyen2023open, deng20213d, kokic2017affordance, iriondo2021affordance, mo2022o2o, kim2014semantic, kim2015interactive].  [kim2014semantic] formulates this task as the point cloud segmentation task from the extracted geometric features.  [kim2015interactive] voxelizes the point cloud and creates affordance maps with interactive manipulation.  [kokic2017affordance, iriondo2021affordance] focuses on the object grasping affordances and some specific applications.  [mo2022o2o] learns the affordance for object-object interactions.  [ngyen2023open] addressed the task of open-vocabulary affordance learning. Overall affordance learning is a very challenging task even the shape of point cloud is given. Instead of predicting the point-wise contact labels, in our method, we propose to make some hypotheses to simplify human-object interactions, making it more tractable without significantly compromising accuracy.
@@ 3.1 Overview
Motion Representations. We denote a 3D HOI sequence as ${\bm{x}}=\{{\bm{x}}^{h},{\bm{x}}^{o}\}$. It consists of human motion sequence ${\bm{x}}^{h}\in\mathbb{R}^{L\times D^{h}}$ and object motion sequence ${\bm{x}}^{o}\in\mathbb{R}^{L\times D^{o}}$, where $L$ denotes the length of the sequence. For ${\bm{x}}^{h}$, we adopt the redundant representation widely used in human motion generation [Guo_2022_CVPR] with $D^{h}=263$, which include pelvis velocity, local joint positions, velocities and rotations of other joints in the pelvis space, and binary foot-ground contact labels. For the object motion sequence ${\bm{x}}^{o}$, we assume the object geometry is given as an input, and thus we only need to estimate its 6DoF poses in the generation, i.e., $D^{o}=6$. We represent each object instance as a point cloud of 512 points ${\bm{p}}\in\mathbb{R}^{512\times 3}$.
@@
Diffusion Model for 3D HOI Generation. Given a prompt ${\bm{c}}=({\bm{d}},{\bm{p}})$, consisting of a textual description ${\bm{d}}$ and the object instance’s point cloud ${\bm{p}}$, a diffusion model $p_{\theta}({{\bm{x}}_{t-1}}|{{\bm{x}}_{t}},{\bm{c}})$***We use superscripts $h$ and $o$ to denote human and object squence, respectively. Without a superscript, it means the 3D HOI sequence, containing both ${\bm{x}}^{h}$ and ${\bm{x}}^{o}$. Subscript is used for the diffusion denoising step. learns the reverse diffusion process to generate clean data from a Gaussian noise ${{\bm{x}}_{t}}$ with $T$ consecutive denoising steps
where $t$ is the denoising step. Following [tevet2023human], instead of predicting the noise $\mathbf{\epsilon}_{t}$ in each diffusion step, our diffusion model $M_{\theta}$ with parameters $\theta$ predicts the final clean motion ${\bm{x}}_{0}=M_{\theta}({{\bm{x}}_{t}},t,{\bm{c}})$. We sample $\mathbf{x}_{t-1}\sim\mathcal{N}({\bm{\mu}}_{t},\Sigma_{t})$ and compute the mean as in [nichol2021improved]
@@
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\beta_{t}\in(0,1)$ are the variance schedule. $\Sigma_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ [Ho2020-ew] is a variance scheduler of choice. Similar to ${\bm{x}}_{t}$, ${\bm{\mu}}_{t}$ consists of ${\bm{\mu}}^{h}_{t}$ and ${\bm{\mu}}^{o}_{t}$, corresponding to human and object motion, respectively.
Figure 3: Illustration of HOI-DM architecture for coarse 3D HOIs generation. It has two branches designed for generating human and object motions individually. A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions.
Simply adopting the diffusion model described in Eq.(1) would impose huge burden to the model, which requires joint generation of human and object motion and more critically, enforcement of their intricate interactions to follow the input textual description. In this paper, we propose HOI-Diff for 3D HOIs generation, which disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of three modules. Fig. 2 shows the overview of our proposed approach. We introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM), which can produce diverse yet consistent motions, capturing the intricate interplay and mutual interactions between humans and objects (Sec. 3.2). To ensure physically plausible contact between humans and objects, we propose a novel affordance prediction diffusion model (APDM) (Sec. 3.3), whose output will be used as classifier guidance (Sec. 3.4) to correct the interactions at each diffusion step of human/object motion generation.
@@ 3.2 Coarse 3D HOIs Generation
First, we introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM) to generate human and object motions that are roughly coherent. As shown in Fig. 3, it consists of two Transformer models [vaswani2017attention], human motion diffusion model (MDM) $M^{h}$ and object MDM $M^{o}$, which work similar to [tevet2023human]. Specifically, at the diffusion step $t$, they take the text description and noisy motions ${\bm{x}}_{t}^{h}$ and ${\bm{x}}_{t}^{o}$ as input and predicts clean human and object motions ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
To enhance the learning of interactions of the human and object when generating their motion, we introduce a Communication Module ($CM$) designed for exchanging feature representations between the human MDM $M^{h}$ and the object MDM $M^{o}$. $CM$ is a Transformer block that receives the intermediate feature ${\bm{f}}^{h},\bm{f}^{o}$ from both $M^{h}$ and $M^{o}$. It then processes these inputs to generate refined updates based on the cross attention mechanism [vaswani2017attention]. The updated feature representations $\tilde{{\bm{f}}_{h}}$ and $\tilde{{\bm{f}}_{o}}$ of the human and object are then conditioned on each other, which are then fed into the subsequent layers of their respective branches to estimate clean human and object motion ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively. The $CM$ is inserted at the 4th transformer layer for human MDM and the last layer for object MDM, which was empirically found to work better.
Given the limited data availability for 3D HOI generation, during training, the human motion model $M^{h}$ finetunes a pretrained human MDM [tevet2023human]. This fine-tuning is critical to ensure the smoothness of the generated human motions. We ablate this design choice in LABEL:ablation.
Figure 4: Illustration of APDM architecture for affordance estimation. Affordance information of human contact labels, object contact positions, and binary object states are represented together as a noise variable, which is fed into the Transformer encoder to generate clean affordance estimation. The object point cloud and textual prompt are taken as conditional input.

### 15 Mar 2024 v2

# https://neu-vi.github.io/HOI-Diff/

@@ Abstract
We address the problem of generating realistic 3D human-object interactions (HOIs) driven by textual prompts. To this end, we take a modular design and decompose the complex task into simpler sub-tasks. We first develop a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioned on the input text, and encourage coherent motions by a cross-attention communication module between the human and object motion generation branches. We also develop an affordance prediction diffusion model (APDM) to predict the contacting area between the human and object during the interactions driven by the textual prompt. The APDM is independent of the results by the HOI-DM and thus can correct potential errors by the latter. Moreover, it stochastically generates the contacting points to diversify the generated motions. Finally, we incorporate the estimated contacting points into the classifier-guidance to achieve accurate and close contact between humans and objects. To train and evaluate our approach, we annotate BEHAVE dataset with text descriptions. Experimental results on BEHAVE and OMOMO demonstrate that our approach produces realistic HOIs with various interactions and different types of objects.
Figure 1: HOI-Diff generates realistic motions for 3D human-object interactions given a text prompt and object geometry. Please see the sup. mat. for video results. Darker color indicates later frames in the sequence. Best viewed in color.
@@ 1 Introduction
The generation of natural and physically plausible 3D HOIs involves humans interacting with dynamic objects in various ways according to the text prompts, thereby posing several challenges. First, the variability of object shapes makes it particularly challenging to generate semantically meaningful contact between the human and object to avoid floating objects. Second, the generated HOIs should be faithful to the input text prompts as there are many plausible interactions between human and the same object (e.g., a person carries a chair, sits on a chair, pushes or pulls a chair). Text-driven 3D HOI synthesis with a diverse set of interactions is still under-explored. Third, the development and evaluation of 3D HOI synthesis models requires a high-quality human motion dataset with various HOIs and textual descriptions, but existing datasets lack either HOIs [11, 34] or textual descriptions [3].
@@
On one hand, recent methods [23, 17, 12, 42, 58, 52, 44, 33] can synthesize realistic human motions for HOIs for static objects only. They usually synthesize the motion in the last mile of interaction, i.e. the motion between the given starting human pose and the final interaction pose, and overlook the movement of the objects when the human is interacting with them. On the other hand, existing methods for motion generation with dynamic objects do not adequately reflect real-world complexity. For instance, they focus on grasping small objects [10], provide the object motion as conditioning [25], predict deterministic interactions between the human and the same object without the diversity [54, 37], consider only a small set of interactions (e.g. sit/lift [23], sit/lie down [12], sit [17, 58, 33], grasp [52, 44]), or investigate a single type of object (e.g. chair [17, 58]).
@@
In this paper, we go beyond and propose HOI-Diff for 3D HOIs synthesis involving humans interacting with different types of objects in diverse ways, which are both physically plausible and semantically faithful to the textual prompt, as shown in Fig. 1. Our key insight is to decompose 3D HOIs synthesis into three modules: (a) coarse 3D HOIs generation that extends the human motion diffusion model [45] to a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioning on the input text prompt. To encourage coherent motions, we develop a cross-attention communication module, exchanging information between the human and object motion generation models; (b) affordance prediction diffusion model (APDM) that estimates the contacting points between the human and object during the interactions driven by the textual prompt. Our APDM does not rely on the results of the HOI-DM and thus can recover from its potential errors. Moreover, it stochastically generates the contacting points to diversity the generated motions; and (c) affordance-guided interaction correction that incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects. Compared with designing a monolithic model, our proposed HOI-Diff disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of the three modules.
@@
To our best knowledge, HOI-Diff is one of the first text-driven approach s capable of generating realistic motions for various HOIs and different types of objects during the interactions. While our approach pioneers this capability, we note the existence of concurrent works such as  [9, 24, 49], which also explore similar directions. For both training and evaluation purposes, we annotate each video sequence in BEHAVE dataset [3] with text descriptions, which mitigates the issue of severe data scarcity for text-driven 3D HOIs generation. In addition, we evaluate our approach on the OMOMO dataset [25], which focuses on the manipulation of two hands. Extensive experiments validate the effectiveness and design choices of our approach, particularly for dynamic objects, thereby enabling a set of new applications in human motion generation.
@@ 2 Related Work
Human Motion Generation with Diffusion Models. The denoising diffusion models have been widely used 2D image generations [39, 40, 36] and achieved impressive results. Recent work [56, 45, 5, 18, 38, 1, 2, 4, 6, 41, 43, 46, 51, 55, 57, 59, 53] apply the diffusion model in the task of human motion generation. While these methods have successfully generated human motion, they usually generate isolated motions in the free space without considering the objects the human is interacting with. Our method is primarily focused on motion generation with human-object interactions.
@@
Scene- and Object-Aware Human Motion Generation. Recent works condition motion synthesis on scene geometry [15, 60, 48, 50]. This facilitates the understanding of human-scene interactions. However, the motion fidelity is compromised due to the lack of paired full scene-motion data. Other approaches [23, 17, 12, 42, 58, 33] instead focus on the interactions with the objects and can produce realistic motions. However, they focus on interacting with static objects with limited interactions. OMOMO [25] proposes a conditional diffusion framework that can generate full-body motion from the object motion. The object motion is needed as input in OMOMO, whereas our method can jointly synthesize human motion and object motion. IMoS [10] synthesizes the full-body human along with the 3D object motions from textual inputs. Although they can generate realistic interactions, IMoS only focuses on grasping small objects with hands. InterDiff [54] predicts whole-body interactions with dynamic objects. However, it focuses on the motion prediction task, where the future motion needs to be forecasted based on past motion information and object information. Note that the interaction type is deterministic, given the object and past motion information. Different from this, we tackle the motion synthesis task, where the interaction with the same object can be controlled by the text prompt. Recently, there has been a surge of interest in the text-driven synthesis of 3D human-object interactions for dynamic objects, resulting in the development of concurrent works [9, 24, 49].
@@
Multi-Human Motion Generation. Recent works [41, 26, 41] incorporate human-to-human interactions into the motion diffusion process. They usually also adopt cross-attention to exchange information from different human motions. Different from them, our goal is to generate coarse HOIs instead of doing so in a single run, which alleviates the burden of the model.
@@
Affordance Estimation. The affordance estimation on 3D point cloud is studied in [30, 7, 22, 16, 29, 20, 21].  [20] formulates this task as the point cloud segmentation task from the extracted geometric features.  [21] voxelizes the point cloud and creates affordance maps with interactive manipulation.  [22, 16] focuses on the object grasping affordances and some specific applications.  [29] learns the affordance for object-object interactions.  [30] addressed the task of open-vocabulary affordance learning. Overall affordance learning is a very challenging task even the shape of point cloud is given. Instead of predicting the point-wise contact labels, in our method, we propose to make some hypotheses to simplify human-object interactions, making it more tractable without significantly compromising accuracy.
@@ 3.1 Overview
Motion Representations. We denote a 3D HOI sequence as ${\bm{x}}=\{{\bm{x}}^{h},{\bm{x}}^{o}\}$. It consists of human motion sequence ${\bm{x}}^{h}\in\mathbb{R}^{L\times D^{h}}$ and object motion sequence ${\bm{x}}^{o}\in\mathbb{R}^{L\times D^{o}}$, where $L$ denotes the length of the sequence. For ${\bm{x}}^{h}$, we adopt the redundant representation widely used in human motion generation [11] with $D^{h}=263$, which include pelvis velocity, local joint positions, velocities and rotations of other joints in the pelvis space, and binary foot-ground contact labels. For the object motion sequence ${\bm{x}}^{o}$, we assume the object geometry is given as an input, and thus we only need to estimate its 6DoF poses in the generation, i.e., $D^{o}=6$. We represent each object instance as a point cloud of 512 points ${\bm{p}}\in\mathbb{R}^{512\times 3}$.
Diffusion Model for 3D HOI Generation. Given a prompt ${\bm{c}}=({\bm{d}},{\bm{p}})$, consisting of a textual description ${\bm{d}}$ and the object instance’s point cloud ${\bm{p}}$, a diffusion model $p_{\theta}({{\bm{x}}_{t-1}}|{{\bm{x}}_{t}},{\bm{c}})$ [^1] We use superscripts $h$ and $o$ to denote human and object sequence, respectively. Without a superscript, it means the 3D HOI sequence, containing both ${\bm{x}}^{h}$ and ${\bm{x}}^{o}$. Subscript is used for the diffusion denoising step. learns the reverse diffusion process to generate clean data from a Gaussian noise ${{\bm{x}}_{t}}$ with $T$ consecutive denoising steps
where $t$ is the denoising step. Following [45], instead of predicting the noise $\mathbf{\epsilon}_{t}$ in each diffusion step, our diffusion model $M_{\theta}$ with parameters $\theta$ predicts the final clean motion ${\bm{x}}_{0}=M_{\theta}({{\bm{x}}_{t}},t,{\bm{c}})$.
@@
We sample $\mathbf{x}_{t-1}\sim\mathcal{N}({\bm{\mu}}_{t},\Sigma_{t})$ and compute the mean as in [31]
@@
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\beta_{t}\in(0,1)$ are the variance schedule. $\Sigma_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ [14] is a variance scheduler of choice. Similar to ${\bm{x}}_{t}$, ${\bm{\mu}}_{t}$ consists of ${\bm{\mu}}^{h}_{t}$ and ${\bm{\mu}}^{o}_{t}$, corresponding to human and object motion, respectively.
Figure 3: Illustration of HOI-DM architecture for coarse 3D HOIs generation. It has two branches designed for generating human and object motions individually. A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions. The human motion model $M^{h}$ finetunes a pretrained human MDM [45].
Figure 3: Illustration of HOI-DM architecture for coarse 3D HOIs generation. It has two branches designed for generating human and object motions individually. A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions. The human motion model $M^{h}$ finetunes a pretrained human MDM [45].
Figure 4: Illustration of APDM architecture for affordance estimation. Affordance information of human contact labels, object contact positions, and binary object states are represented together as a noise variable, which is fed into the Transformer encoder to generate clean estimation. The object point cloud and textual prompt are taken as conditional input.
Simply adopting the diffusion model described in Eq.(1) would impose a huge burden on the model, which requires joint generation of human and object motion and more critically, enforcement of their intricate interactions to follow the input textual description. In this paper, we propose HOI-Diff for 3D HOIs generation, disentangling motion generation for humans and objects and estimation of their contacting points. They are later integrated to form coherent and diverse HOIs, which reduces the complexity and burden for each of the three modules, leading to better generation performance as evidenced by our experiments.
Fig. 2 shows the overview of our proposed approach. We introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM), which can produce diverse yet consistent motions, capturing the intricate interplay and mutual interactions between humans and objects (Sec. 3.2). To ensure physically plausible contact between humans and objects, we propose a novel affordance prediction diffusion model (APDM) (Sec. 3.3), whose output will be used as classifier guidance (Sec. 3.4) to correct the interactions at each diffusion step of human/object motion generation.
@@ 3.2 Coarse 3D HOIs Generation
First, we introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM) to generate human and object motions that are roughly coherent. As shown in Fig. 4, it consists of two Transformer models [47], human motion diffusion model (MDM) $M^{h}$ and object MDM $M^{o}$, which work similar to [45]. Specifically, at the diffusion step $t$, they take the text description and noisy motions ${\bm{x}}_{t}^{h}$ and ${\bm{x}}_{t}^{o}$ as input and predict clean human and object motions ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
To enhance the learning of interactions of the human and object when generating their motion, we introduce a Communication Module ($CM$) designed for exchanging feature representations between the human MDM $M^{h}$ and the object MDM $M^{o}$. $CM$ is a Transformer block that receives the intermediate feature ${\bm{f}}^{h},\bm{f}^{o}$ from both $M^{h}$ and $M^{o}$. It then processes these inputs to generate refined updates based on the cross attention mechanism [47]. The updated feature representations $\tilde{{\bm{f}}_{h}}$ and $\tilde{{\bm{f}}_{o}}$ of the human and object are then conditioned on each other, which are then fed into the subsequent layers of their respective branches to estimate clean human and object motion ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively. The $CM$ is inserted at the 4th transformer layer for human MDM and the last layer for object MDM, which was empirically found to work better.
Given the limited data availability for 3D HOI generation, during training, the human motion model $M^{h}$ finetunes a pretrained human MDM [45]. This fine-tuning is critical to ensure the smoothness of the generated human motions. We ablate this design choice in Sec. 4.3. Object MDM is trained from scratch. We modify the input and output linear layers to take in the object motion which has a different dimension from the human motion. More methodology details of HOI-DM can be found in the supplementary material.